A firm's average revenue function AR=−18−7,5Q+Q2.AR=−18−7,5Q+Q2.



Determine the number of units to be produced and sold to maximise revenue.

1
Expert's answer
2022-04-05T09:49:55-0400

"AR=\frac{R}{Q}\Rightarrow R=QAR=-18Q-7.5Q^2+Q^3\rightarrow \max \\Since\\\underset{Q\rightarrow +\infty}{\lim}\left( -18Q-7.5Q+Q^3 \right) =+\infty ,\\there\,\,is\,\,no\,\,\max value. \\"